OpenStudy (anonymous):

a=14.2, b=9.3, C=9.2 Find the area of triangle ABC

OpenStudy (anonymous):

a=14.2, b=9.3, C=9.2°

OpenStudy (anonymous):

Do you think you can draw a picture? Because the only reason that a triangle would have A, B, and C listed as a measurement would be for a^2 + b^2 = c^2

OpenStudy (jdoe0001):

@tigers49 have you done the Law of Cosines yet?

OpenStudy (anonymous):

There's no picture. All i know is i have to use 1/2(a)(b) sin C. After that I have no idea what to do.

OpenStudy (anonymous):

@jdoe0001 would I use that formula?

OpenStudy (jdoe0001):

well, that's what I'd use to find "c"

OpenStudy (jdoe0001):

but if you haven't covered it, then you may not be doing trig, so, it may not apply

OpenStudy (anonymous):

I have went over it. I'm just confused the answer I got was 10.55

OpenStudy (jdoe0001):

so, the triangle will look like 

OpenStudy (anonymous):

The other part I was confused on was finding the remaining sides and the angles of the triangle of ABC

OpenStudy (jdoe0001):

finding the other angles is just using the Law of Sines as far as find the Area just with the sides ONLY, I'd use Heron's formula but if you haven't covered those, not sure if this would apply

OpenStudy (anonymous):

Okay so in that case would I have to use K=1/2(a)(b) sin C first. Then use Heron's formual to find the area of the triangle?

OpenStudy (anonymous):

I have went over heron's formula and the law of Sines and Cosines. I'm just lost on when and how to use them